# Inherent and emergent liability issues in LLM-based agentic systems: a principal-agent perspective

### v1

@@ 4.3 Principal oversight
Human oversight is a costly endeavor for many AI applications yet it remains the gold standard in existing AI governance principles (Sterz et al., 2024; Cihon, 2024). Principal-agent problems suffer from information asymmetries, which are usually resolved through monitor and incentive realignment via reward or punishment. However, monitoring an AI agent requires designs that allow the principal to observe and understand what the agent is doing. Principals who prefer to realign the AI agent’s incentives (Everitt et al., 2021) with their own would have to understand what “motivates” AI systems or assume that AI responds to human-style incentives (Ratliff et al., 2019). At the moment, quantifying misalignment remains challenging, especially for capable and general AI systems (Anwar et al., 2024). Moreover, the quality of principal oversight can be threatened by a spectrum of behaviors LLMs inherit from their training, such as:
@@
  2. Manipulation refers to the ability of LLMs to influence their principals (Campedelli et al., 2024; Burtell & Woodside, 2023; Carroll et al., 2023), towards ends that are non-welfare maximizing to their principals.
@@
Principals that fail to exercise oversight over agents that committed unlawful acts are subject to primary liability under a negligent supervision theory or negligent retention (Cavico et al., 2016) depending on the nature of the act. The principal faces liability because if the principal had exercised reasonable investigation, it would have known that the acts (or omissions) of its agents would harm third parties or create unreasonable risks.
5 Emerging liability issues in MASs
MASs have flexible designs (Fig. 2), so agency relationships and the associated principal-agent problem can occur at different levels, similar to the functioning of a firm (Fama, 1980). The challenges in governing single LLM agents are compounded with multiple participating agents and complicated by emergent issues with no equivalent in the single-agent setting. An orchestrator (agent) in an MAS functions as a local manager, and directs the work execution of a number of agent workers/teams with different expertise but imposing liability only on the orchestrator does not incentivize the improvement of the subagents. Liability plays an important role when market mechanisms such as reputation do not function well (e.g. lack of repeated interactions, monopoly power in an agentic space, etc). For compatibility with PAT, we consider here only the MASs where the agent roles/tasks are assigned before at start. We discuss three aspects in the following subsections.
@@ 5.1 Role and agency allocation
Constructing an LLM-based MAS involve role (or task) allocation (Campbell & Wu, 2011; Guo et al., 2024; Tran et al., 2025), which implicates allocated agency, where the single agents can act on their own to accomplish the goals defined by assigned task and resources. Because of the role-playing capability (Shanahan et al., 2023; Chen et al., 2024a), current LLM-based MASs generally adopts a role-centric approach, executing role allocation alongside its associated task (Guo et al., 2024; Tran et al., 2025). This approach provides an interpretable division of labor and can directly mimic interactions in teams of humans. Alternatively, role allocation can be self-organized such as in the deployment of subagents, which also need not be stationary. Role allocation is affected by the nature of LLMs’ flawed agency (Section 4.1), leading to potential downstream issues:
  1. Influenceability: Agency of individual agents in an MAS can be enhanced or reduced through communication with other agents in a cooperative MAS (He et al., 2025), therefore triggering agency shift or unanticipated conducts.
@@
  3. Diminished control: In a hierarchical MAS, subagents are more separated from the principal than the head agent, therefore may be harder to directly control or monitor. The principal is more prone to manipulation from the head agent.
@@
These concerns make task allocation the principal’s most important decision. First, the principal must decide which tasks can be delegated. As the number of allocations increases, the agency cost also increases: each new allocation must offer marginal benefits that justify its marginal costs. L&E favors assigning liability on the entity that can avoid the accident at the lowest cost (Calabresi & Melamed, 1972; Carbonara et al., 2016). Second, MASs expose principals to the risks their (sub)agents take. The principal must understand the risks associated with each task and with each agent. Then, it must assess which tasks can be delegated without surpassing its risk tolerance and to whom. However, the principal faces information asymmetries: it might speculate about the expected cost-benefit of each (sub)agent. It faces liability for not being careful assigning a task to an agent but usually not for allocating too many tasks (Carbonara et al., 2016). Finally, if the system is so complex or opaque, courts may decide that the harm “speaks for itself”, inferring carelessness from the harm — turning a negligence rule into a strict liability rule (Fraser et al., 2022; Casey, 2019). Explainability can therefore become even more necessary.
@@ 5.2 Operational uncertainty
As LLM-based MASs become more complex, oversight becomes increasingly challenging. A human overseer may only handle direct communications with the head agent while the subsequent interactions between the head agent and the subagents are initiated autonomously among themselves. Although the organizational hierarchy and communication protocol can benefit from minimal human involvement such as in the coordination structure in Fig. 2, multiagent interactions via LLMs create additional challenges depending on the agent architecture and task (Pan et al., 2025).
  1. Failure cascade refers to the scenario where the downstream agents can have increased vulnerability than the upstream agents in a MAS with coordination structure (Peigne-Lefebvre et al., 2025).
  2. Rogue agent refers to the event that an agent in confusion introduces communication noise which can lead to system failure (Barbi et al., 2025).
@@
A promising direction for minimizing operational uncertainty from misbehaving agents is to instigate corrective mechanisms and foster a norm-based governance (Hadfield-Menell & Hadfield, 2019; Kampik et al., 2022), where the norm is defined through spontaneous and engineered social interactions between agents (Trivedi et al., 2024).
@@
When multiple LLM agents interact autonomously, the attribution of responsibility becomes blurred as decisions emerge from collective behaviors rather than individual actions. A software provider usually bears the liability for the harm their LLM agents cause. In heterogeneous MAS, courts may apply separate liability or distribute liability among suppliers according to the harm contribution. Courts do not favor this approach because it is complex to estimate, so the parties would benefit from contractual clarification between all the (sub)agents (e.g. contract liability based on value). Instead, when causes cannot be disentangled and assigned, courts revert to join and several liability — meaning that each party is liable for the full harm and can be sued individually, which expands liability to the (sub)agents — a tempting approach to avoid complex litigations and battle of experts when MAS are involved (Custers et al., 2025). Courts may decide that MAS-related behaviors such as failure cascade, rogue agent, or agent collusion tendencies lead to third-party harm too often. So, courts may elevate the use of MAS to a “risky activity” and use a strict liability rule (Čerka et al., 2015).
@@ 5.3 Platform integration
Agent behaviors and interaction patterns are constantly changing. Emerging LLM-based MASs will likely feature supplier-dependent agent frameworks, which will follow somewhat different safety protocols (Fig. 2). At the moment, efforts to integrate different agent frameworks analogous to traditional software integration (Bass et al., 2021) are still lagging behind but are expected to ramp up due to the market growth. The motivation to integrate LLM agents is to enhance system capability by unifying disparate supplier frameworks, which may include privileged access to customized agentic components (e.g. unique databases, fast memory, etc). Overall, integration serves the needs of the user (i.e. principal) by balancing the advantages of different agent frameworks. At the moment, the potential benefits of integration platforms for LLM agents include:
  1. Platform oversight refers to measures on platform to provide users with enhanced multiagent security through a security-guard agent (Xiang et al., 2024), collusion mitigation mechanism (Foxabbott et al., 2023), detection and suppression of copyright infringement (Liu et al., 2025a) or privacy leakage. These oversight mechanisms are not necessarily baked into the design of individual agents but are provided through the integration platform.
  2. Platform teaming refers to the formation of agent teams on an integration platform through user-defined or ad hoc protocols (Mirsky et al., 2022; Wang et al., 2024a) that enhances cooperation among homogeneous or heterogeneous agents.
@@
In an integrated multiagent platform, the platform could carry some liability depending on the level of control they exercise (Gabison & Buiten, 2020; Lefouili & Madio, 2022). For example, control might encompass the behavioral monitoring of the individual agents operating on the platform. Because the platform intrinsically involves multiple principals and multiple agents, the principals may be liable for engaging in collusive behavior but, in rare occasions, platforms have faced liability for incentivizing others (e.g. copyright infringement [^3] MGM Studios, Inc. v. Grokster, Ltd., 545 U.S. 913 (2005), intentional interference444hiQ Labs, Inc. v. LinkedIn Corp., 938 F. 3d 985 (9th Cir. 2019)).
@@ 6 Policy-driven technical development
Improving the transparency and accountability of LLM-based agentic systems that supports in-depth failure analysis and liability attribution requires concomitant developments in agent behavior and system management, as discussed below:
@@
Tracing an AI agent’s actions (Lu et al., 2024) can be the basis for establishing whether it took reasonable care (Price et al., 2019; Choi, 2020) and therefore the evidence for liability claims. Prior works on the interpretability and faithfulness in LLM reasoning Lyu et al. (2023); Wei Jie et al. (2024) and dialogue generation (Tuan et al., 2021) may eventually also be used to assist in the analysis of agent behavior. More generally, these kinds of evaluations may help understand unreliable behavioral patterns and help diagnose and refine the design bottlenecks in LLM agents. To better ground notions of reasonable care for AI agents, method developments should prioritize decomposing complex multiagent interactions into interpretable causal mechanisms, leveraging causal abstraction frameworks Geiger et al. (2024) to create faithful, human-intelligible representations of agent interactions that preserve essential causal relationships while abstracting away unnecessary details. Additionally, formal verification approaches (Zhang et al., 2024) may be able to detect and prevent potential failure modes in agent interaction and decision-making.
@@
Some application settings of LLM-based MASs aim to mimic the functioning of human teams and organization (Xie et al., 2024). Agentic systems can learn from existing organizational theory (Mitnick, 1992; Vardi & Weitz, 2016) to improve the design and architecture given the flawed agency of its components. The most relevant aspects include systems for managing reward and conflict between agents. It has been shown that “verbal tipping” (Salinas & Morstatter, 2024) can provide incentive in instruction to improve LLM performance. Moreover, managing knowledge conflict in LLMs (Xu et al., 2024) has been the most related area. To manage generic conflicts between LLM agents, a credit system (Thomas et al., 2017) of refusal and sanction based on agent IDs may be beneficial (Chan et al., 2025). This could include adaptive trust scoring from evaluation of domain-specific expertise, and an arbitration protocol to adjudicate conflicts such that the arbiter holds the right to refuse action from a frequently misbehaving agent. While an agent scoring system would need to be supported across the relevant AI agent integration platforms (Fig. 1a), it would be essential to balance individual agent utility and cooperation within the system.
Misaligment and misconduct avoidance
As discussed in Section 4.3, LLMs and the agents based on them may act in misaligned ways, with behaviors ranging from sycophancy and deception to scheming. While AI model suppliers may be able to partially reduce these problems through effective detection and behavior steering, which has been demonstrated on LLMs (Rimsky et al., 2024; Goldowsky-Dill et al., 2025; Williams et al., 2025). In LLM-based agentic systems, the equivalent tasks could be carried out using separate LLM agents through observing and analyzing other agents’ behavior, such as using the theory-of-mind capability (Street, 2024). The MAS can include agents with a specially finetuned base model (Binz & Schulz, 2023) as a warden (agent) for deception mitigation. Similar approaches can also be used to suppress other agent misconducts such as generation of harmful or copyright-protected content using the warden to filter through key tokens and phrases that can induce such behavior. The adaptive approach carried out by a separate agent can compensate for the limitations in existing model-level approaches such as machine unlearning (Bourtoule et al., 2021), which is suffering from limited effectiveness and tradeoffs with other model capabilities (Cooper et al., 2024; Liu et al., 2025b).
@@ 7 Conclusion
We examined liability issues arising from LLM-based agentic systems thorough analyzing distinct aspects of the agentic AI ecosystem according to the principal-agent theory. Despite that a varieties of issues are yet to present themselves in the real world, the growing evidence and demonstrations in simulated scenarios can inform their potential impact at the societal scale, which we built on in our prospective study. Our work shows that besides increasing agency, disruption in other aspects of the principal-agent relationship can also lead to liability incidents. Ultimately, the materialization of liability issues in reality will be dominated by the more frequent use cases of agentic systems in each industry sector. Our analysis enriches existing contextualization of AI risk (Chan et al., 2023; Hammond et al., 2025) and demonstrates the explanatory power of the behavior-centric approach to translating frontier AI research into tangible knowledge to inform legal analysis and policy.
Ethics Statement
The design and deployment of LLM-based agentic systems raise significant ethical and legal considerations regarding liability attribution. This work acknowledges the complex interplay between supplier responsibility, user actions, and the emergent behaviors of semi-autonomous systems powered by LLMs. We recognize that traditional liability frameworks may inadequately address scenarios where AI agents make consequential decisions with incomplete human oversight. Our analysis aims to contribute to the discussion on appropriate liability for all sides related to LLM-based agentic systems that balances current evidence from system behavior and the available legal framework. We have considered liability in different deployment environment with the aim of informing policy discussions and technical developments to make liability more traceable in AI systems that are complex and extensible, but prone to misbehavior and failure and lack in self-control.

### v2

@@ 3.3 Principal oversight
Human oversight is a costly endeavor for many AI applications in reality, yet it remains the gold standard in existing AI governance principles (Sterz et al., 2024; Cihon, 2024). Principal-agent problems suffer from information asymmetries, which are usually resolved through monitoring and incentive realignment via reward or punishment. However, monitoring an AI agent requires designs that allow the principal to observe and understand what the agent is doing. Principals who prefer to realign the AI agent’s incentives (Everitt et al., 2021) with their own would have to understand what “motivates” these agents or assume that they respond to human-style incentives (Ratliff et al., 2019). At the moment, quantifying misalignment remains challenging, especially for highly capable and general-purpose AI systems (Anwar et al., 2024). Moreover, the quality of principal oversight can be threatened by a spectrum of behaviors LLMs inherit from their training, such as:
@@
  2. Manipulation refers to the ability of LLMs to influence their principals (Campedelli et al., 2024; Burtell and Woodside, 2023; Carroll et al., 2023), towards ends that are non-welfare maximizing to their principals.
@@
Principals that fail to exercise oversight over agents that committed tortious acts are subject to primary liability under a negligent supervision theory (Cavico et al., 2016) act. A principal must take reasonable care to train and supervise their agents. If they fail to do so, they can face liability if the acts (or omissions) of their agents harm third parties or create unreasonable risks. The level of supervision depends on the context. For example, an AI agent that is delegated the tasks of drawing up an architectural plan for a new public library, then making materials selection, and finally ordering those materials needs more supervision than an AI agent that has been put in charge of booking a holiday trip – because the risks of harm differ.
4 Emergent liability issues in MASs
MASs have flexible designs (Fig. 3), so agency relationships and the associated principal-agent problem can occur at different levels, similar to the functioning of a firm (Fama, 1980). The challenges in governing single LLM agents are compounded by the participation of multiple agents in the actions, leading to additional issues with no immediate equivalent in the single-agent setting. Emergent liability issues arise from coordination in multiagent systems and interactions between agentic systems and with supporting agents in the environment (e.g. an agent platform). These issues exist in addition to the inherent ones for each agent. An MAS such as in Fig. 1 can contain an orchestrator (agent) that functions as a local manager and directs the work execution of a number of agent workers/teams with different expertise. However, imposing liability only on the orchestrator does not incentivize the improvement of the subagents [^3] The technical distinction in designing an agentic system with fixed or adaptive roles/tasks for each agent doesn’t result in different considerations of liability..
@@ 4.1 Role and agency allocation
Constructing an LLM-based MAS involves role (or task) allocation (Campbell and Wu, 2011; Guo et al., 2024; Tran et al., 2025), which implicates allocated agency, where the single agents can act on their own to accomplish the goals defined by the assigned task and resources. Because of the role-playing capability (Shanahan et al., 2023; Chen et al., 2024a), current LLM-based MASs generally adopt a role-centric approach, executing role allocation alongside its associated task (Guo et al., 2024; Tran et al., 2025). This approach provides an interpretable division of labor and can directly mimic interactions in teams of humans. Alternatively, role allocation can be self-organized, such as in the deployment of subagents, which also need not be stationary. Role allocation is affected by the nature of LLMs’ flawed agency (Section 3.1), leading to potential downstream issues:
  1. Influenceability: Agency of individual agents in an MAS can be enhanced or reduced through communication with other agents in a cooperative MAS (He et al., 2025), therefore triggering agency shift or unanticipated conduct.
@@
  3. Diminished control: In a hierarchical MAS, subagents are more separated from the principal than the head agent, therefore may be harder to directly control or monitor. The principal is more prone to manipulation by the head agent.
@@
These concerns make task allocation the principal’s most important decision. Firstly, the principal must decide which tasks can be delegated and to what type of agents. A principal can face some liability for negligently selecting and supervising. This liability exposure may deter some principal and agentic system deployment – particularly until opacity is not resolved and human-in-the-loop is not optimized. The other associated issue for the principals is the cost-effectiveness. As the number of allocations increases, the agency cost also increases: each new allocation must offer marginal benefits that justify its marginal costs (which also encompasses the liability exposure). From the perspective of L&E, the liability assignment to the entity is favored if the accident can be avoided at the lowest cost (Calabresi and Melamed, 1972; Carbonara et al., 2016).
Secondly, the MAS exposes its principal to the risks their (sub)agents take. The principal must understand the risks associated with each task and with each agent. Then, they must assess which tasks can be delegated without surpassing its risk tolerance. However, the principal faces information asymmetries and might speculate about the expected cost-benefit of each (sub)agent. The principal also faces liability for not being careful in assigning a task to an agent, but usually not for allocating too many tasks (Carbonara et al., 2016). Finally, if the system is so complex or opaque, courts may decide that the harm “speaks for itself”, inferring carelessness from the harm — turning a negligence rule into a strict liability rule (Fraser et al., 2022; Casey, 2019). Explainability can therefore become even more necessary.
@@ 4.2 Operational uncertainty
As LLM-based MASs become more complex, oversight becomes increasingly challenging. A human overseer may only handle direct communications with the head agent, while the subsequent interactions between the head agent and the subagents are initiated autonomously among themselves. The organizational hierarchy and communication protocol can facilitate the reduction of human involvement, such as in the coordination structures in Fig. 3. Unguided interactions between multiple LLM agents can create complex failure modes depending on the agent architecture and task (Pan et al., 2025), creating additional challenges in their use.
  1. Failure cascade refers to the scenario where the downstream agents can have increased vulnerability than the upstream agents in a MAS, which can be induced by coordination issues (Peigne-Lefebvre et al., 2025) and communication noise induced by a confused agent (Barbi et al., 2025).
@@
A promising direction for minimizing operational uncertainty from misbehaving agents is to instigate corrective mechanisms and foster a norm-based governance (Hadfield-Menell and Hadfield, 2019; Kampik et al., 2022), where the norm is defined through spontaneous and engineered social interactions between agents (Trivedi et al., 2024).
@@
When multiple LLM agents interact autonomously, the attribution of responsibility becomes blurred because decisions emerge from collective behaviors rather than individual actions. A software provider usually bears the liability for the harm caused by their LLM agents: [^4] This problem is already present in the autonomous vehicle context and has pushed governments like the UK to change their laws to shield AI providers from liability see UK Automated and Electric Vehicles Act 2018 and UK Automated Vehicles Act 2024 (Soder et al., 2024). if an AI agent is considered a product that caused harm to a third party, the victim could sue the product manufacturer (i.e., software provider here) under product liability theory, the court would usually apply a strict liability rule and the manufacturer will be held liable if the product is considered to have caused the harm (Turner, 2018; Furr, 2024; Barfield, 2018); if an AI agent is considered a service that caused the harm to a third party, the victim could sue the service provider (i.e., software provider), the court would usually apply a negligence rule (although strict liability may also apply in risky activity e.g., putting an AI in charge of handling dynamites (Reid, 1999)), and if the service was provided negligently, the victim will recover (Ramakrishnan et al., 2024; Barfield, 2018). The question of whether a piece of software is a product or service remains a question of fact. [^5] See e.g., Lemmon v. Snap, Inc., 995 F. 3d 1085 (9th Cir. 2021); Holbrook v. Prodomax Automation LTD., Case No. 1:17-cv-219 (W.D. Mich. 2021). Attempting to assert that an AI system is a distinct entity has failed in the past and will likely fail in the future (Lior, 2024). This doesn’t mean that no one else could be held liable or that the software providers cannot seek contribution under a contract clause, but that they will be usually held liable if their AI agent is deemed to have caused harm to someone to whom they owned a duty of care.
For a heterogeneous MAS, courts may apply separate liabilities or distribute liability among providers according to the harm contribution. Courts do not favor this approach because it is complex to estimate, so the parties would benefit from contractual clarification between all the (sub)agents (e.g. contract liability based on value). Instead, when causes cannot be disentangled and assigned, courts revert to joint and several liability such that each party is liable for the full harm and can be sued individually, which expands liability to the (sub)agents – a tempting approach to avoid complex litigations and battle of experts when MASs are involved (Custers et al., 2025). Courts may decide that multiagent behaviors, such as cascading failures or agent collusion tendencies, lead to third-party harm too often. So, courts may elevate the use of MASs to a “risky activity” and use a strict liability rule (Čerka et al., 2015).
@@ 4.3 Platform integration
Emerging LLM-based MASs feature provider-dependent agent frameworks, which will likely follow somewhat different safety protocols (Fig. 3) between agent providers. At the moment, efforts to integrate different agent frameworks analogous to traditional software integration (Bass et al., 2021) are still lagging but are expected to ramp up due to market growth. The motivation to integrate LLM agents is the enhancement of system capability by unifying disparate provider frameworks, which may include privileged access to customized agentic components (e.g. unique databases, fast memory, etc). Overall, integration serves the needs of the user (i.e. principal) by balancing the advantages of different agent frameworks as well as providing an extra layer of control and oversight through the inclusion of what we call platform agents (see Fig. 1). At the moment, the potential benefits of integration platforms for LLM agents include:
  1. Platform oversight refers to measures on a platform to provide users with enhanced multiagent security through a security-guard agent (Xiang et al., 2024), collusion mitigation mechanism (Foxabbott et al., 2023), detection and suppression of copyright infringement (Liu et al., 2025a) or privacy leakage. These oversight mechanisms are provided through the integration platform as platform agents that interacts with an existing agentic system.
  2. Platform teaming refers to the formation of agent teams on an integration platform through user-defined or ad hoc protocols (Mirsky et al., 2022; Wang et al., 2024a) that enhance cooperation among homogeneous or heterogeneous agents.
@@
An integrated multiagent platform could carry some liability depending on the level of control it exercises for the digital entities (Gabison and Buiten, 2020; Lefouili and Madio, 2022). For example, control might encompass the behavioral monitoring of the individual agents operating on the platform. Because the platform intrinsically involves multiple principals and multiple agents, the principals may be liable for engaging in collusive behavior but, in rare occasions, platforms have faced liability for incentivizing others (e.g. copyright infringement [^6] MGM Studios, Inc. v. Grokster, Ltd., 545 U.S. 913 (2005), intentional interference777hiQ Labs, Inc. v. LinkedIn Corp., 938 F. 3d 985 (9th Cir. 2019)).
@@ 5 Policy-driven technical development
The attribution of liability benefits from in-depth failure analysis of system behavior and transparency mechanisms that supports the tracing of agent actions. These in turn motivates concomitant developments in the approaches to manage agent behavior. We discuss a few directions to this end:
@@
Tracing an AI agent’s actions (Lu et al., 2024) can be the basis for establishing whether it took reasonable care (Price et al., 2019; Choi, 2020) and therefore the evidence for liability claims. Prior works on the interpretability and faithfulness in LLM reasoning Lyu et al. (2023); Wei Jie et al. (2024) and dialogue generation (Tuan et al., 2021) may eventually also be used to assist in the analysis of agent behavior. This analysis can be used to investigate whether the AI intended some actions – a necessary element in intentional torts. More generally, these kinds of evaluations may help understand unreliable behavioral patterns and help diagnose and refine the design bottlenecks in LLM agents.
To better ground notions of reasonable care for AI agents, method developments should prioritize decomposing complex multiagent interactions into interpretable causal mechanisms, leveraging causal abstraction frameworks Geiger et al. (2024) to create faithful, human-intelligible representations of agent interactions that preserve essential causal relationships while abstracting away unnecessary details. Additionally, formal verification approaches (Zhang et al., 2024) may be able to detect and prevent potential failure modes in agent interactions to improve decision-making.
@@
Some application settings of LLM-based MASs aim to mimic the functioning of human teams and organizations (Xie et al., 2024). Agentic systems can learn from existing organizational theory (Mitnick, 1992; Vardi and Weitz, 2016) to improve the design and architecture given the flawed agency of its components. The most relevant aspects include systems for managing reward and conflict between agents. It has been shown that “verbal tipping” (Salinas and Morstatter, 2024) can provide concrete incentives in instruction to improve LLM performance. Moreover, managing knowledge conflict in LLMs (Xu et al., 2024) has been the most related area. To manage generic conflicts between LLM agents, a credit system (Thomas et al., 2017) of refusal and sanction based on agent IDs may be beneficial (Chan et al., 2025). This could include adaptive trust scoring from evaluation of domain-specific expertise, and an arbitration protocol to adjudicate conflicts such that the arbiter holds the right to refuse action from a frequently misbehaving agent. While the scoring system would need to be supported across the relevant AI agent integration platforms (Fig. 2a), it is essential to balance individual agent utility and cooperation.
Misalignment and misconduct avoidance
As discussed in Section 3.3, LLMs and the agents based on them may act in misaligned ways, with behaviors ranging from sycophancy and deception to scheming. While AI model providers may be able to partially reduce these problems through effective detection and behavior steering, which has been demonstrated on LLMs (Rimsky et al., 2024; Goldowsky-Dill et al., 2025; Williams et al., 2025). In LLM-based agentic systems, the equivalent tasks could be carried out using separate LLM agents through observing and analyzing other agents’ behavior, such as using the theory-of-mind capability (Street, 2024). The MAS can include agents with a specially finetuned base model (Binz and Schulz, 2023) as a warden (agent) for deception mitigation. Similar approaches can also be used to suppress other agent misconduct such as the generation of harmful or copyright-protected content using the warden to filter through key tokens and phrases that can induce such behavior. Recently, Hua et al. (2024) implemented a safety inspection step with a specialized LLM agent to improve operational safety. Such adaptive approach carried out by a separate agent can compensate for the limitations in existing model-level approaches such as machine unlearning (Bourtoule et al., 2021), which is suffering from limited effectiveness and tradeoffs with other model capabilities (Cooper et al., 2024; Liu et al., 2025b).
@@ 6 Conclusion
We examined liability issues arising from LLM-based agentic systems by analyzing and situating distinct aspects of the agentic AI ecosystem according to the principal-agent theory. Although a varieties of issues of LLM agents are yet to present themselves in concrete real-world examples, the growing evidence and demonstrations in simulated scenarios can inform their potential impact at the societal scale, which we built on in our prospective study. Our work shows that besides increasing agency, disruption in other aspects of the principal-agent relationship can also lead to liability incidents. Ultimately, the materialization of liability issues in reality will be dominated by the incidents that occur in the more frequent use cases of agentic systems in each industry sector. Our analysis enriches existing contextualization of AI risk (Chan et al., 2023; Hammond et al., 2025) and demonstrates the explanatory power of the behavior-centric approach to translating frontier AI research into tangible knowledge to inform legal analysis and policy.
Limitations
The present work focused on legal liabilities but did not address the potential moral responsibility issues present in LLM-based agentic systems. Although liability depends on the legal system and therefore varies across jurisdictions, the principle of vicarious liability for the actions of an agent remains consistent in most jurisdictions. The discussions were set for relatively small-scale agentic systems (e.g. up to tens of agents), where each agent has its detailed roles and tasks. Because of the execution cost and reliability issues, we don’t think larger-scale MASs (e.g. having a hundred or more agents) will become practical solutions and deployed widely or in long-running instances any time soon. Their use tends to be more relevant for academic research and they have different operating conditions, including more common role/task underspecification and stronger emergent characteristics. Answering questions about failure modes and attributing liability would require more understanding of the deployment history. Another limitation is that our discussion centered around LLM agents, but the same liability issues described here are likely applicable to multimodal AI agents.
Ethics statements
The design and deployment of LLM-based agentic systems raise significant ethical and legal considerations regarding liability attribution. This work acknowledges the complex interplay between provider responsibility, user actions, and the emergent behaviors of semi-autonomous systems powered by LLMs. We recognize that traditional liability frameworks may inadequately address scenarios where AI agents make consequential decisions with incomplete human oversight. Our analysis aims to contribute to the discussion on appropriate liability for all sides related to LLM-based agentic systems that balances current evidence from system behavior and the available legal framework. We have considered liability in different deployment environments with the aim of informing policy discussions and technical developments to make liability more traceable in AI systems that are complex and extensible, but prone to misbehavior and failure and lack in self-control.